In a regular 13-gon, each interior angle measures____degrees, and each exterior angle measures___degrees. Round to the nearest tenth if necessary.

Find the sum of the measures of the interior angles of a decagon.

Answer by Alan3354(30993)

In a regular 13-gon, each interior angle measures____degrees, and each exterior angle measures___degrees. Round to the nearest tenth if necessary.
------------
The sum of the exterior angles is 360º for ALL polygons.
Each exterior angle = 360/13 =~ 27.7º
Each interior angle = 180 - 27.7 = 152.3º
----------------------
Find the sum of the measures of the interior angles of a decagon.
Sum = (n-2)*180 = 8*180
= 1440º